# No Compromise in Solution Quality: Speeding Up Belief-dependent Continuous POMDPs via Adaptive Multilevel Simplification

### v1

@@ 4.2 Resimplification strategy: LAZY-Gap
The PT resimplification strategy from previous section assure that no overlap is present (Fig. 3) at each non-leaf posterior belief and we know the optimal action to take. However, it can inflict a redundant computational burden. We can handle the overlap only at the root of the belief tree and use the bounds over optimal value function according to (42) and (43). Since we already presented the resimplification strategy based on the simplification levels, our second resimplification strategy will be based on the distance between reward bounds. However, the bounds (42) and (43) can be utlizied also with resimplifcation strategy based on simplification levels directly. Yet, this is out of the scope of this paper. In this section we present a lazy variant of resimplifcation strategy. We recite again that no necessarily lazy variant shall be tied to the gap. In LAZY variant, the overlap is checked solely at the root $b_{k}$ of the whole belief tree. In this approach three scenarios can be encountered at each belief node.
@@
  • In the presence of overlap at the root $b_{k}$ (Eq. (38) is not satisfied), we shall prune actions according to (39) and Alg. 2 and commence resimplification routine for the non pruned actions based on the resimplification strategy.
@@ 5 Adaptive Simplification in the Setting of MCTS
In the previous sections, we described the application of the adaptive simplification paradigm when the belief tree is given or its construction is not coupled with the solution. We now turn to anytime setting where the belief tree is not given. Instead, the belief tree construction is coupled with the estimation of the action-value function (19) at each belief action node. Such an approach is commonly used in Monte Carlo tree search (MCTS) methods based on an exploration strategy, e.g. Upper Confidence Bound (UCB) as in (15). Our goal is to suggest a resimplifcation strategy so that exactly the same belief tree as without simplification would be constructed. Also the same optimal action is identified with and without simplification. To support general belief-dependent rewards we select PFT-DPW as the baseline, as mentioned in Section 1.1.
@@ 5.2 Guaranteed Belief Tree Consistency
Since the simplification paradigm substituted UCB (15) by the bounds (47) and (48) the belief tree construction is coupled with these quantities as opposed to the situation with the given belief tree, we now shall address this question. If there is an overlap between bounds on UCB for different actions, we can no longer guarantee the same belief tree will be constructed with and without simplification.
@@
Consider two algorithms, constructing a belief tree. Assume every common sampling operation for the two algorithms uses the same seed. The two algorithms are tree consistent if two belief trees constructed by the algorithms are identical in terms of actions, observations, and visitation counts.
Our approach relies on a specific procedure for selecting actions within the tree. Since each simulation the MCTS descends down the tree with a single return lace as in (19), on the way down it requires the action maximizing UCB (15) we shall eliminate overlap at each belief node as described in section 3.4. Further we restate the action selection procedure described in section 3.4 with particular action dependent constant from eq.  (37) and (38) rendering the UCB bounds from (47) and (48).
Our action selection is encapsulated by Alg. 8, which is different from the procedure used in PFT-DPW. On top of DPW as in (Sunberg and Kochenderfer, 2018) with parameters $k_{a}$ and $\alpha_{a}$, instead of selecting an action maximizing the UCB (15), at every belief node we mark as a candidate action the one that maximizes the lower bound $\underline{\mathrm{UCB}}$ as such
@@
If $\forall a\neq\tilde{a}$, $\underline{\mathrm{UCB}}(h\tilde{a})\geq\overline{\mathrm{UCB}}(ha)$, there is no overlap (Fig. 6 (c)) and we can declare that $\tilde{a}$ is identical to $a^{\ast}$, i.e., the action that would be returned by PFT using (15) and the tree consistency has not been affected. Otherwise, the bounds must be tightened, so ensure the tree consistency. We examine the $ha$ siblings of $h\tilde{a}$, which satisfy $a\neq\tilde{a}:\underline{\mathrm{UCB}}(h\tilde{a})<\overline{\mathrm{UCB}}(ha)$ (Fig. 6 (a)). Our next step is to tighten the bounds by resimplification (Fig. 6 (b)) until there is no overlap using the valid resimplification strategy according to Definition 1. Note that here we cannot use the “lazy variant” from section 4.2 due to fact that the MCTS requires to select an action going down to the tree, see line $12$ of the Algorithm 7.
@@ 5.3 A Detailed Algorithm Description
Now we introduce our efficient variant of the Particle Filter Tree (PFT) presented in (Sunberg and Kochenderfer, 2018). We call our approach Simplified Information-Theoretic Particle Filter Tree (SITH-PFT). SITH-PFT (Alg. 7) incorporates the adaptive simplification into PFT-DPW. We adhere to the conventional notations as in (Sunberg and Kochenderfer, 2018) and denote by $G_{\text{PF}(m)}(bao)$ a generative model receiving as input the belief $b$, an action $a$ and an observation $o$, and producing the posterior belief $b^{\prime}$. For belief update, we use a particle filter based on $n_{x}$ state samples. A remarkable property of our efficient variant is the consistency of the belief tree. In other words, PFT and SITH-PFT have same belief tree constructed with (15), while SITH-PFT enjoys substantial acceleration. By $C(ha)$ we denote the set of the children (posterior beliefs corresponding to the myopic observations) of the belief action node uniquely indexed by the history $h$ with concatenated action $a$. The line 13 in Alg. 7 is the DPW technique from (Sunberg and Kochenderfer, 2018) with parameters $k_{o}$ and $\alpha_{o}$. The $N(\cdot)$ is the visitation count of belief or belief action nodes. In MCTS, the $Q$ estimate is assembled by averaging the laces of the returns over simulations see Eq. 19. Each simulation yields a sum of discounted cumulative rewards. Therefore, by replacing the reward with adaptive lightweights bounds (21), we get corresponding discounted cumulative upper and lower bounds over the returns. Averaging the simulations (Alg. 7 lines 28-29), yields the bounds over the action-value function and the UCB bounds used in the routine ActionSelection() to be explained in the next paragraph.
@@ 5.4 Specific Resimplification Strategy: PT-Gap
In this section, we explain the resimplification procedure in more detail. In particular we present specific resimplification strategy and further show that this strategy is valid according to the Definition 1. When some sibling belief action nodes have overlapping bounds (Fig. 3, Fig. 6), we strive to avoid tightening them all at once since fewer resimplifications lead to greater acceleration (speedup). Thus, we choose a single $ha$-node that causes the largest “gap”, denoted by $G$, between its bounds (see Alg. 8 lines 24-30), where $G$ is defined by (46).
@@ 5.5 Guarantees
In this section we first show that resimplification strategy suggested in the previous section is valid.
@@
The resimplification strategy presented in section 5.4 promotes the simplification level of at least one reward in the rollout or belief tree. Alternatively, all the rewards are at maximal simplification level $n_{\mathrm{max}}$. In other words the suggested resimplifcation strategy is valid.
We provide the complete proof in Appendix 10.2. Having proved the validity of suggested resimplification strategy, we proceed to the monotonicity and convergence of UCB bounds from (47) and (48).
@@
The UCB bounds are monotonic as the function of the number of resimplifciations and after at most $n_{\mathrm{max}}\cdot M$ resimplifications we have that
@@
We placed the proof in Appendix 10.3. Now, using Lemma 2, we prove that SITH-PFT (Alg. 7) yields same belief tree and same best action as PFT.
@@
We provide full proof of Theorems 2 and 3 in Appendix 10.4 and 10.5 respectively. We showed that for any valid resimplification strategy SITH-PFT is guaranteed to construct same belief tree as PFT and select the same best action at the root. From Lemma 1, our resimplification strategy is valid. Thus, we achieved the desired result.
@@ 6.1 Novel Bounds Over Differential Entropy Estimator
As one of our key contributions, we now derive novel analytical bounds for the differential entropy estimator from (Boers et al., 2010). These bounds can then be used within our general simplification framework presented in the previous sections.
To calculate differential entropy
$\mathcal{H}(b(x_{k}))=-\int b(x_{k})\cdot\log\left(b(x_{k})\right)\mathrm{d}x_{k},$
one must have access to the manifold representing the belief. In nonparametric setting this manifold is out of reach. We have to resort to approximations. Several approaches exist. One of them is using Kernel Density Estimation (KDE) as done, e.g., by Fischer and Tas (2020). Here, however, we consider the method proposed by Boers et al. (2010). This method builds on top of usage of motion and observation models such that
@@
Utilizing the chosen simplification (53), we now introduce the following upper and lower bounds on 54.
@@
See proof in Sec. 10.6. The Thm. 4 accommodates different sets $A^{s}_{k}\neq A^{s}_{k+1}$. These sets as explained above denote sets of particle indices from $b_{k}$ and $b_{k+1}$ for simplification level $s$. In general each of these sets can have its own simplification level. However, this is out of the scope of this paper. Here, both sets $A^{s}_{k}$, $A^{s}_{k+1}$ have same simplification level as well as the number of levels. Yet, the number of particles at each level can vary between $A^{s}_{k}$ and $A^{s}_{k+1}$. Each subsequent level (low to high) defines a larger set of indices such that higher levels of simplification (i.e. more samples) correspond to tighter and lower levels of simplification correspond to looser bounds. Note that the bounds (56) and (57) actually use the original and simplified beliefs so it settles with Eq. 20 and 21.
Importantly, by caching the shared calculations of the both bounds in same time instance, we never repeat the calculation of these values and obtain maximal speedup.
@@
Without compromising on the solution’s quality we are accelerating the online decision making process.
@@ 6.2 Bounds Properties and Analysis
We now turn to analysis of the bound and investigation of their properties. Allow us to start from computational complexity. We then examine monotonicity and convergence of the bounds and reuse of calculations.
@@ 6.2.1 Computational complexity
The(56) and (57) suggest that the bounds are cheaper to calculate than $\hat{\mathcal{H}}$ from (54), with complexity of $O(n_{x}^{s}\cdot n_{x})$ instead of $O(n_{x}^{2})$, where $n_{x}^{s}\triangleq|A^{s}_{k}|\equiv|A^{s}_{k+1}|$. Altogether, time saved for all belief nodes in the tree will result in the total speedup of our approach.
@@ 6.2.2 Monotonicy and Convergence
See proof in Sec. 10.7. Finally, bounding (54) using Theorem 4 corresponds, in our general framework from Section 3, to (20).
@@ 6.2.3 Re-use of Calculations
We provide the theoretical time complexity analysis using the specific bounds (from Section 6.1) in the Appendix 10.8. Now we are keen to present our simulations.
@@ 7 Simulations and Results
First, we study empirically the specific simplification and bounds from Section 6 and show that they become tighter as the number of particles increases. We, then benchmark our algorithms for planning in the setting of a given belief tree (Section 4) and in an anytime MCTS setting (Section 5). In the former setting, we compare SITH-BSP and LAZY-BSP against Sparse Sampling (Kearns et al., 2002). In an anytime MCTS setting, we compare SITH-PFT with PFT-DPW (Sunberg and Kochenderfer, 2018). This performance evaluation is conducted considering two problems, as discussed next.
@@ 7.1 Problems under Consideration
We proceed to the description of the evaluated problems. In both problems the immediate reward for $b^{\prime}$ is
@@ 7.1.1 Continuous Light Dark
$\mathbb{P}_{T}(x^{\prime}|x,a)=\mathcal{N}(x+a,I\cdot\sigma_{T}),$
@@
Figure 8: The plot shows the evolution of belief in terms of sets of particles along the actual trajectory of the robot. The color of the particles from yellow to red illustrate the evolution of the belief over time. The green ellipses represent the parametric Gaussian belief covariances obtained from update by Kalman filter. The canvas color here is $\sigma_{\mathcal{O}}=\sigma_{T}=0.075$. (a) Our first scenario. (b) Our second scenario.

### v2

@@ 4.2 Resimplification strategy: LAZY-Gap
The PT resimplification strategy from previous section assure that no overlap is present (Fig. 3) at each non-leaf posterior belief and we know the optimal action to take. However, it can inflict a redundant computational burden. We can handle the overlap only at the root of the belief tree and use the bounds over optimal value function according to (43) and (44). Since we already presented the resimplification strategy based on the simplification levels, our second resimplification strategy will be based on the distance between reward bounds. However, the bounds (43) and (44) can be utilized directly also with the resimplifcation strategy based on simplification levels. Yet, this is out of the scope of this paper.
In this section we present a lazy variant of the resimplifcation strategy. In a LAZY variant, the overlap is checked solely at the root $b_{k}$ of the whole belief tree. In this approach three scenarios can be encountered at each belief node.
@@
  • In the presence of an overlap at the root $b_{k}$ (Eq. (39) is not satisfied), we shall prune actions according to (40) and Alg. 2 and commence a resimplification routine for the non pruned actions based on the resimplification strategy.
@@ 5 Adaptive Simplification in the Setting of MCTS
In the previous sections, we described the application of the adaptive simplification paradigm when the belief tree is given or its construction is not coupled with the solution. We now turn to an anytime setting where the belief tree is not given. Instead, the belief tree construction is coupled with the estimation of the action-value function (20) at each belief action node. Such an approach is commonly used in Monte Carlo tree search (MCTS) methods based on an exploration strategy, e.g. Upper Confidence Bound (UCB) as in (16). Our goal is to suggest a resimplifcation strategy so that exactly the same belief tree as without simplification would be constructed. Also the same optimal action is identified with and without simplification. To support general belief-dependent rewards we select PFT-DPW as the baseline, as mentioned in Section 1.1.
@@ 5.2 Guaranteed Belief Tree Consistency
Since the simplification paradigm substituted UCB (16) by the bounds (48) and (49), the belief tree construction is coupled with these quantities, as opposed to the situation with the given belief tree. If there is an overlap between bounds on UCB for different actions, we can no longer guarantee the same belief tree will be constructed with and without simplification.
@@
Imagine two algorithms, constructing a belief tree. Assume every common sampling operation for the two algorithms uses the same seed. The two algorithms are tree consistent if two belief trees constructed by the algorithms are identical in terms of actions, observations, and visitation counts.
Our approach relies on a specific procedure for selecting actions within the tree. Since in each simulation the MCTS descends down the tree with a single return lace as in (20), on the way down it requires the action maximizing UCB (16) we shall eliminate overlap at each belief node as described in section 3.4. Further we restate the action selection procedure described in section 3.4 with particular action dependent constant from eq.  (38) and (39) rendering the UCB bounds from (48) and (49).
Our action selection is encapsulated by Alg. 8, which is different from the procedure used in PFT-DPW. On top of DPW as in Sunberg and Kochenderfer (2018) with parameters $k_{a}$ and $\alpha_{a}$, instead of selecting an action maximizing the UCB (16), at every belief node we mark as a candidate action the one that maximizes the lower bound $\underline{\mathrm{UCB}}$ as such
@@
If $\forall a\neq\tilde{a}$, $\underline{\mathrm{UCB}}(h\tilde{a})\geq\overline{\mathrm{UCB}}(ha)$, there is no overlap (Fig. 7 (c)) and we can declare that $\tilde{a}$ is identical to $a^{\ast}$, i.e., the action that would be returned by PFT using (16) and the tree consistency has not been affected. Otherwise, the bounds must be tightened, so ensure the tree consistency. We examine the $ha$ siblings of $h\tilde{a}$, which satisfy $a\neq\tilde{a}:\underline{\mathrm{UCB}}(h\tilde{a})<\overline{\mathrm{UCB}}(ha)$ (Fig. 7 (a)). Our next step is to tighten the bounds by resimplification (Fig. 7 (b)) until there is no overlap using the valid resimplification strategy according to Definition 1.
Remark: Note that here we cannot use the “lazy variant” from Section 4.2 due to the fact that the MCTS requires selecting an action going down to the tree, see line $12$ of Algorithm 7. Therefore, if the UCB bounds do still overlap, we cannot assure that the same acton will be selected as in case of UCB itself.
@@ 5.3 A Detailed Algorithm Description
Now we introduce our efficient variant of the Particle Filter Tree (PFT) presented in Sunberg and Kochenderfer (2018). We call our approach Simplified Information-Theoretic Particle Filter Tree (SITH-PFT). SITH-PFT (Alg. 7) incorporates the adaptive simplification into PFT-DPW. We adhere to the conventional notations as in Sunberg and Kochenderfer (2018) and denote by $G_{\text{PF}(m)}(bao)$ a generative model receiving as input the belief $b$, an action $a$ and an observation $o$, and producing the posterior belief $b^{\prime}$. For belief update, we use a particle filter based on $n_{x}$ state samples. A remarkable property of our efficient variant is the consistency of the belief tree. In other words, PFT and SITH-PFT have the same belief tree constructed with (16), while SITH-PFT enjoys substantial acceleration. By $C(ha)$ we denote the set of the children (posterior beliefs corresponding to the myopic observations) of the belief action node uniquely indexed by the history $h$ with concatenated action $a$. Line 13 in Alg. 7 is the DPW technique from Sunberg and Kochenderfer (2018) with parameters $k_{o}$ and $\alpha_{o}$. The $N(\cdot)$ is the visitation count of belief or belief action nodes. In MCTS, the $Q$ estimate is assembled by averaging the laces of the returns over simulations see Eq. 20. Each simulation yields a sum of discounted cumulative rewards. Therefore, by replacing the reward with adaptive lightweights bounds (22), we get corresponding discounted cumulative upper and lower bounds over the returns. Averaging the simulations (Alg. 7 lines 28-29), yields the bounds over the action-value function and the UCB bounds used in the routine ActionSelection() to be explained in the next paragraph.
@@ 5.4 Specific Resimplification Strategy: PT-Gap
In this section, we explain the resimplification procedure in more detail. In particular we present a specific resimplification strategy and further show that this strategy is valid according to Definition 1. When some sibling belief action nodes have overlapping bounds (Fig. 3, Fig. 7), we strive to avoid tightening them all at once since fewer resimplifications lead to greater acceleration (speedup). Thus, we choose a single $ha$-node that causes the largest “gap”, denoted by $G$, between its bounds (see Alg. 8 lines 24-30), where $G$ is defined by (47).
@@ 5.5 Guarantees
In this section we first show that the resimplification strategy suggested in the previous section is valid.
@@
The resimplification strategy presented in Section 5.4 promotes the simplification level of at least one reward in the rollout or belief tree. Alternatively, all the rewards are at the maximal simplification level $n_{\mathrm{max}}$. In other words the suggested resimplifcation strategy is valid.
We provide the complete proof in Appendix 11.2. Having proved the validity of the suggested resimplification strategy, we proceed to the monotonicity and convergence of UCB bounds from (48) and (49).
@@
The UCB bounds are monotonic as a function of the number of resimplifications and after at most $n_{\mathrm{max}}\cdot M$ resimplifications we have that
@@
We provide the proof in Appendix 11.3. Now, using Lemma 2, we prove that SITH-PFT (Alg. 7) yields the same belief tree and the same best action as PFT.
@@
We provide the full proofs of Theorems 2 and 3 in Appendix 11.4 and 11.5, respectively. We showed that for any valid resimplification strategy SITH-PFT is guaranteed to construct the same belief tree as PFT and select the same best action at the root. From Lemma 1, our resimplification strategy is valid. Thus, we achieved the desired result.
@@ 6.1 Novel Bounds Over Differential Entropy Estimator
As one of our key contributions, we now derive novel analytical bounds for the differential entropy estimator from Boers et al. (2010). These bounds can then be used within our general simplification framework presented in the previous sections. To calculate differential entropy
$\mathcal{H}(b(x_{k}))\triangleq-\int b(x_{k})\cdot\log\left(b(x_{k})\right)\mathrm{d}x_{k},$
one must have access to the manifold representing the belief. In a nonparametric setting this manifold is out of reach. We have to resort to approximations. Several approaches exist. One of them is using Kernel Density Estimation (KDE) as done, e.g., by Fischer and Tas (2020). Here, however, we consider the method proposed by Boers et al. (2010). This method builds on top of usage of motion and observation models such that
@@
Utilizing the chosen simplification (54), we now introduce the following upper and lower bounds on (55).
@@
See proof in Appendix 11.6. Theorem 4 accommodates different sets $A^{s}_{k}\neq A^{s}_{k+1}$. These sets denote sets of particle indices from $b_{k}$ and $b_{k+1}$ for simplification level $s$. In general, each of these sets can have its own simplification level. However, this is out of the scope of this paper. Here, both sets $A^{s}_{k}$, $A^{s}_{k+1}$ have the same simplification level, as well as the number of levels. Yet, the number of particles at each level can vary between $A^{s}_{k}$ and $A^{s}_{k+1}$. Each subsequent level (low to high) defines a larger set of indices such that higher levels of simplification (i.e. more samples) correspond to tighter and lower levels of simplification correspond to looser bounds. Note that the bounds (57) and (58) actually use the original and simplified beliefs so it settles with Eqs. (21) and (22).
Importantly, by caching the shared calculations of both bounds in the same time instance, we never repeat the calculation of these values and obtain maximal speedup. Without compromising on the solution’s quality we are accelerating the online decision making process.
@@ 6.2 Bounds Properties and Analysis
We now turn to analysis of the bounds and investigation of their properties. Allow us to start from computational complexity. We then examine monotonicity and convergence of the bounds and reuse of calculations.
@@ 6.2.1 Computational complexity
Eqs. (57) and (58) suggest that the bounds are cheaper to calculate than $\hat{\mathcal{H}}$ from (55), with complexity of $O(n_{x}^{s}\cdot n_{x})$ instead of $O(n_{x}^{2})$, where $n_{x}^{s}\triangleq|A^{s}_{k}|\equiv|A^{s}_{k+1}|$. Altogether, time saved for all belief nodes in the tree will result in the total speedup of our approach.
@@ 6.2.2 Monotonicy and Convergence
See proof in Appendix 11.7. Finally, bounding (55) using Theorem 4 corresponds, in our general framework from Section 3, to (21).
@@ 6.2.3 Re-use of Calculations
We provide the theoretical time complexity analysis using the specific bounds (from Section 6.1) in Appendix 11.8. Now we are keen to present our simulations.
7 Adaptation Overhead
Whereas the bounds presented in Section 6 are incremental repeated resimplifications may lead to actually slower decision-making. This overhead is caused by additional algorithmics introduced by the resimplification routine. We can anticipate such scenarios when the actions are symmetrical in terms of the reward. However, as we empirically observed and will shortly present in the next section, in the setting of given belief tree the cases where the simplification is beneficial prevail. Especially in the LAZY variant since there the Alg. 5 engages resimplification routine only at the root of the belief tree.
In the setting of MCTS the situation is slightly more complicated. In UCB we cannot prune actions for eternity but only dismiss up until the next arrival to the belief node. This is because when MAB (defined in Section 2.3.3) converges it switches the current best action with arrivals to the belief node; such a behavior necessitates our simplification approach to tighten the bounds for many candidate actions. As a result in a MCTS setting we obtain less speedup than in the setting of a given belief tree considering LAZY variant (Alg. 5). Nevertheless in some problems the simplification approach is invaluable, as for example, in the problem described in Section 8.1.3 and investigated in Section 8.3.5. Importantly, we can further accelerate resimplification routines by parallelization. However, this is out of the scope of this paper. All our implementations are single threaded.
@@ 8 Simulations and Results
First, we study empirically the specific simplification and bounds from Section 6 and show that they become tighter as the number of particles increases. We, then benchmark our algorithms for planning in the setting of a given belief tree (Section 4) and in an anytime MCTS setting (Section 5). In the former setting, we compare SITH-BSP and LAZY-BSP against Sparse Sampling (Kearns et al., 2002). In an anytime MCTS setting, we compare SITH-PFT with PFT-DPW (Sunberg and Kochenderfer, 2018) and IPFT (Fischer and Tas, 2020). This performance evaluation is conducted considering three problems, as discussed next.
@@ 8.1 Problems under Consideration
We proceed to the description of the evaluated problems. In two first problems the immediate reward for $b^{\prime}$ is
@@ 8.1.1 Continuous Light Dark
$\mathbb{P}_{T}(x^{\prime}|x,a)=\mathcal{N}(x+a,I\cdot\sigma_{T}),$ (60)
@@
Figure 9: The plot shows the evolution of belief in terms of sets of particles along the actual trajectory of the robot. The color of the particles from yellow to red illustrate the evolution of the belief over time. The green ellipses represent the parametric Gaussian belief covariances obtained from update by Kalman filter. The canvas color here is $\sigma_{\mathcal{O}}=\sigma_{T}=0.075$ as in equations (60) and (61) respectively. (a) Our first scenario. (b) Our second scenario.
@@ 8.1.2 Target Tracking
8.1.3 Safe Autonomous Localization
Our third problem is a variation of the problem presented in Section 8.1.1. Here we change the reward to be the combination of localization reward and safety reward (Zhitnikov and Indelman, 2022a)
$\begin{gathered}\rho(b,a,z^{\prime},b^{\prime})=\overbrace{-\hat{\mathcal{H}}(b,a,z^{\prime},b^{\prime})}^{\text{localization reward}}+\\ +\underbrace{s\Big{(}2\cdot\mathbf{1}_{\{\mathrm{P}(\{x^{\prime}\in\mathcal{X}^{\mathrm{safe},\prime}\}|b^{\prime})\geq\delta\}}(b^{\prime})-1\Big{)}}_{\text{safety reward}}.\end{gathered}$ (63)
Such a safety reward divides the candidate actions into two sets, the safe set and the unsafe. If the safety parameter $s$ is sufficiently large to assure that safe action is selected, these two sets are detached enough in terms of safety reward and the unsafe set is substantially inferior such that there is no point to calculate localization reward precisely over this set of actions. There, we can, without any harm for decision-making outcome, substitute differential entropy by the bounds at the low simplification levels. This aspect makes the simplification paradigm invaluable.